OpenStudy (anonymous):

what are the rules of ln? Like when it asks to subtract, and you have to divide. what are the rules of ln? Like when it asks to subtract, and you have to divide. @Mathematics

OpenStudy (anonymous):

The common algebraic rules of logarithms of any base, including base e, are \[ \ln(ab) = \ln(a) + \ln(b)\] \[ \ln(\frac{a}{b}) = \ln(a) - \ln(b) \] \[ \ln(a^b) = b*\ln(a) \]

OpenStudy (anonymous):

thanx